undeadknight26 (undeadknight26):

How would i solve this for y? 2x + 3y = 1500

OpenStudy (amistre64):

subtract then divide ...

undeadknight26 (undeadknight26):

subtract: 2x = 3y =1500 -2 -2 ???

undeadknight26 (undeadknight26):

@amistre64

OpenStudy (amistre64):

2x is a single term ... you have to treat it as one number to subtract it subtract 2x mutliplication and division can be used to split a term apart; divide 3y by 3

OpenStudy (amistre64):

2x + 3y = 1500 -2x -2x ---------------- 3y = -2x + 1500 /3 /3 ----------------- y = (-2x+1500)/3